solve the initial value problem using power series
(x^2 +2)y"-xy'+y=0
Y(0)=7 ,y'(0)=9
1
Expert's answer
2020-08-26T16:46:05-0400

"(x^2+2)y''-xy'+y=0\newline
y=C_0x^\sigma+C_1x^{\sigma+1}+C_2x^{\sigma+2}+...\\
(x^2+2)*(C_0\sigma(\sigma-1)x^{\sigma-2}+C_1(\sigma+1)\sigma x^{\sigma-1}+...)-x*(C_0\sigma x^{\sigma-1}+C_1(\sigma+1)x^{\sigma}+...)+C_0x^\sigma+C_1x^{\sigma+1}+C_2x^{\sigma+2}+...=0\\
Let\space coefficients \space near \space x^k(where \space k=\sigma-2, \ldots,\infin)\space be\, equal \space to \space zero. \space \\Then \space we \space have \space the \space next\space system:\\
\begin{cases}
2C_0\sigma(\sigma-1)=0\\
2C_1\sigma(\sigma+1)=0\\
C_0\sigma(\sigma-1)+(\sigma+2)(\sigma+1)C_2-C_0\sigma+C_0=0\\
........\\
C_k(\sigma+k-1)^2+2C_{k+2}(\sigma+k+2)(\sigma+k+1)=0
\end{cases}\\
Let\space\sigma=0\space then:\\
C_k=\dfrac{(-1)^{k/2}C_0}{2^{k/2}k!}, \space if \space k \space is\,divisible \space by\, 2, and\space C_k=\dfrac{(-1)^{(k-1)/2}C_1}{2^{(k-1)/2}k!}\space \space otherwise\space .\\
y=\underset{k=0,\space k=k+2}{\sum}\dfrac{(-1)^{k/2}C_0}{2^{k/2}k!}x^k +\underset{k=1,\space k=k+2}{\sum}\dfrac{(-1)^{(k-1)/2}C_1}{2^{(k-1)/2}k!}x^k\newline
y(0)=7=>C_0=7\\
y'(0)=9=>C_1=9\\
y=7\underset{k=0,\space k=k+2}{\sum}\dfrac{(-1)^{k/2}}{2^{k/2}k!}x^k +9\underset{k=1,\space k=k+2}{\sum}\dfrac{(-1)^{(k-1)/2}}{2^{(k-1)/2}k!}x^k\newline"